OpenStudy (anonymous):

How do you write a prime number multiplication expression for 24?

OpenStudy (solomonzelman):

\[2 \times 2 \times 2 \times 3\]

OpenStudy (anonymous):

Figure out the numbers that need to be multiplied together to get 24. Prime numbers are those that are only evenly divisible by themselves (or by 1): 1, 2, 3, 5, 7, 11, 13...

OpenStudy (isaiah.feynman):